# Enhancing Memory and Imagination Consistency in Diffusion-based World Models via Linear-Time Sequence Modeling

### v1

@@ 4 Motivational Experiments
(a) Qualitative results in MemoryS7 for DIAMOND and ours.
(b) Quantitative results in MemoryS7/S9: DIAMOND & ours.
Figure 1: Motivational examples for both qualitative and quantitative evidences to demonstrate that DIAMOND face difficulties in imagining accurate future under long horizon memorization tasks.
To substantiate the memory limitations of the DIAMOND model, we conducted experiments using the MiniGrid MemoryS7 and MemoryS9 environments (Chevalier-Boisvert et al., 2023). These experiments evaluate memory consistency and temporal prediction capabilities in world models. Fig. 1 presents qualitative comparisons among the DIAMOND model, our proposed EDELINE method, and the Oracle world model. The world models were trained on MiniGrid MemoryS7, and their state predictions were evaluated. The upper half of Fig. 1 (a) shows an initial action sequence of four consecutive left turns followed by four no-op actions. DIAMOND and EDELINE then autoregressively predicted the subsequent states. The lower half of Fig. 1 (a) reveals the prediction outcomes. At timestep 9, DIAMOND generated incorrect predictions with an extra key object in the state. The model’s predictions at timesteps 10 to 11 exhibited inaccuracies in the outer wall representations. These prediction errors result from the environment’s partial observability and DIAMOND’s four-frame context limitation. The input of four idle frames at timestep 9 led to loss of earlier state context. In contrast, EDELINE maintained accurate predictions throughout the sequence. This improved performance stems from EDELINE’s architectural design, which addresses the memory constraints inherent in DIAMOND. The accurate prediction at timestep 12 by DIAMOND can be attributed to its access to the 8th state, which enabled accurate predictions at timestep 13 through the correct 12th state context. Fig. 1 (b) illustrates DIAMOND’s performance limitations in both MemoryS7 and MemoryS9 environments due to insufficient memory capacity in partially observable scenarios. In contrast, EDELINE shows near-optimal performance under these conditions.
@@ 5 Methodology
Figure 2: Framework Overview of EDELINE.
The model integrates three principal components: (1) An U-Net-like Next-Frame Predictor enhanced by adaptive group normalization and cross-attention mechanisms, (2) A Recurrent Embedding Module built on Mamba architecture for temporal sequence processing, and (3) A Reward/Termination Predictor implemented through linear layers. The EDELINE framework uses shared hidden representations across the components for efficient world model learning.
Conventional diffusion-based world models (Alonso et al., 2024) demonstrate promise in learning environment dynamics yet face fundamental limitations in memory capacity and horizon prediction consistency. To address these challenges, this paper presents EDELINE, as illustrated in Fig 2, a unified architecture that integrates state space models (SSMs) with diffusion-based world models. EDELINE’s core innovation lies in its integration of SSMs for encoding sequential observations and actions into hidden embeddings, which a diffusion model then processes for future frame prediction. This hybrid design maintains temporal consistency while generating high-quality visual predictions. A Convolutional Neural Network based actor processes these predicted frames to determine actions, thus enabling autoregressive generation of imagined trajectories for policy optimization.
@@ 5.1.1 Recurrent Embedding Module
While DIAMOND, the current state-of-the-art in diffusion-based world models, relies on a fixed context window of four previous observations and actions sequence, the proposed EDELINE architecture advances beyond this limitation through a recurrent architecture for extended temporal sequence processing. Specifically, we provide theoretical evidence in Theorem F.1 to support that recurrent embedding can preserve more information compared with stacked frames inputs. At each timestep $t$, the Recurrent Embedding Module processes the current observation-action pair $(o_{t},a_{t})$ to update a hidden state $h_{t}=f_{\phi}(h_{t-1},o_{t},a_{t})$. The implementation of REM utilizes Mamba (Gu & Dao, 2024b), an SSM architecture that offers distinct advantages for world modeling. This architectural selection is motivated by the limitations of current sequence processing methods in deep learning. Self-attention-based Transformer architectures, despite their strong modeling capabilities, suffer from quadratic computational complexity which impairs efficiency. Traditional recurrent architectures including Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Chung et al., 2014) experience gradient instability issues that affect dependency learning. In contrast, SSMs provide an effective alternative through linear-time sequence processing coupled with robust memorization capabilities via their state-space formulation. The adoption of Mamba emerges as a promising choice due to its demonstrated effectiveness in modeling temporal patterns across various sequence modeling tasks. Section 6.3 presents a comprehensive ablation study that evaluates different architectural choices for the REM.
@@ 5.1.2 Next-Frame Predictor
While motivated by DIAMOND’s success in diffusion-based world modeling, EDELINE introduces significant architectural innovations in its Next-Frame Predictor to enhance temporal consistency and feature integration. At time step $t$, the model conditions on both the last $L$ frames and the hidden embedding $h_{t}$ from the Recurrent Embedding Module to predict the next frame $\hat{o}_{t+1}$. The predictive distribution $p_{\phi}(o^{0}_{t+1}|h_{t})$ is implemented through a denoising diffusion process, where $D_{\phi}$ functions as the denoising network. Let $y_{t}^{\tau}=(\tau,o^{0}_{t-L+1},...,o^{0}_{t},h_{t})$ represent the conditioning information, where $\tau$ represents the diffusion time. The denoising process can be formulated as $o^{0}_{t+1}=D_{\phi}(o^{\tau}_{t+1},y_{t}^{\tau}).$ To effectively integrate both visual and hidden information, $D_{\phi}$ employs two complementary conditioning mechanisms. First, the architecture incorporates (Zheng et al., 2020) layers within each residual block to condition normalization parameters on the hidden embedding $h_{t}$ and diffusion time $\tau$, which establishes context-aware feature normalization (Zheng et al., 2020). This design significantly extends DIAMOND’s implementation, which limits AGN conditioning to $\tau$ and action embeddings only. The second key innovation introduces cross-attention blocks inspired by Latent Diffusion Models (LDMs), which utilize $h_{t}$ and $\tau$ as context vectors. The UNet’s feature maps generate the query, while $h_{t}$ and $\tau$ project to keys and values. This novel attention mechanism, which is absent in DIAMOND, facilitates the fusion of spatial-temporal features with abstract dynamics encoded in $h_{t}$. The observation modeling loss $\mathcal{L}_{\text{obs}}(\phi)$ is defined based on Eq. (45), and can be formulated as follows:
@@ 5.1.3 Reward / Termination Predictor
#Superhuman (↑) 0 N/A 1 8 10 10 9 7 11 13
Mean (↑) 0.000 1.000 0.332 0.956 1.046 1.266 1.124 0.989 1.459 1.866
Median (↑) 0.000 1.000 0.134 0.505 0.289 0.580 0.485 0.270 0.373 0.817
IQM (↑) 0.000 1.000 0.130 0.459 0.501 0.636 0.487 - 0.641 0.940
Optimality Gap (↓) 1.000 0.000 0.729 0.513 0.512 0.433 0.510 - 0.480 0.387
5.1.4 EDELINE World Model Training
The world model integrates an innovative end-to-end training strategy with a self-supervised approach. EDELINE extends the harmonization technique from HarmonyDream (Ma et al., 2024) through the adoption of harmonizers $w_{o}$ and $w_{r}$, which dynamically balance the observation modeling loss $\mathcal{L}_{\text{obs}}(\phi)$ and reward modeling loss $\mathcal{L}_{\text{rew}}(\phi)$. This adaptive mechanism results in the total loss function $\mathcal{L}(\phi)$:
$\begin{split}\mathcal{L}(\phi)=w_{0}\mathcal{L}_{\text{obs}}(\phi)&+w_{r}\mathcal{L}_{\text{rew}}(\phi)+\mathcal{L}_{\text{end}}(\phi)\\ &+\log(w_{o}^{-1})+\log(w_{r}^{-1})\end{split}$ (4)
To optimize computational efficiency while ensuring robust learning, the Next-Frame Predictor learns to utilize hidden embeddings from any timestep through strategic random sampling for $\mathcal{L}_{\text{obs}}(\phi)$. For a sequence of length $T$, it follows:
$\mathcal{L}_{\text{obs}}(\phi)=\|\hat{o}^{0}_{t+1}-o_{t+1}\|^{2},$ (5)
where $i\sim\text{Uniform}\{1,2,\ldots,T-1\},\quad\hat{o}^{0}_{t+1}\sim p(\hat{o}^{0}_{t+1}\mid h_{t}).$ For reward and termination prediction, EDELINE utilizes cross-entropy losses averaged over the sequence, which can be formulated as: $\mathcal{L}_{\text{rew}}(\phi)=\frac{1}{T}\sum_{t=1}^{T}\text{CrossEnt}(\hat{r}_{t},r_{t}),$ $\mathcal{L}_{\text{end}}(\phi)=\frac{1}{T}\sum_{t=1}^{T}\text{CrossEnt}(\hat{d}_{t},d_{t}).$ This unified training approach, combining random sampling strategies with dynamic loss harmonization, demonstrates superior efficiency compared to DIAMOND’s separate network methodology, as validated in our results presented in Section 6. Moreover, the quantitative analysis presented in Appendix G reveals substantial reductions in world model training duration.
5.2 Agent Behavior Learning
To enable fair comparison and demonstrate the effectiveness of EDELINE’s world model architecture, the agent architecture adopts the same optimization framework as DIAMOND. Specifically, the agent integrates policy $\pi_{\theta}$ and value $V_{\theta}$ networks with REINFORCE value baseline and Bellman error optimization using $\lambda$-returns (Alonso et al., 2024). The training framework executes a procedure with three key phases: experience collection, world model updates, and policy optimization. This method, as formalized in Algorithm 1, follows the established paradigms in model-based RL literature (Łukasz Kaiser et al., 2020; Hafner et al., 2020; Micheli et al., 2023; Alonso et al., 2024). To ensure reproducibility, we provide extensive details in the Appendix, with documentation of objective functions, and the hyperparameter configurations in Appendices J, I, respectively.
6 Experiments
This section presents our experimental results of EDELINE. We provide details of our experimental setup in Appendix E.
6.1 Atari 100k Experiments
@@
Following standard evaluation paradigms for world models, we evaluate EDELINE on the Atari 100k benchmark. For performance quantification, we adopt the human-normalized score (HNS) (Wang et al., 2016), which measures agent performance relative to human and random baselines:
@@
normed score=agent scorerandom scorehuman scorerandom $\text{normed score}=\frac{\text{agent score}-\text{random score}}{\text{human score}-\text{random score}}$ (6)
Figure 3: Mean, median and interquartile HNS.
@@
Fig. 3 presents stratified bootstrap confidence intervals following Agarwal et al. (2021)’s recommendations for point estimate limitations. It can be observed that EDELINE exhibits exceptional performance across this benchmark. Our approach surpasses human players in 13 games with a superhuman mean HNS of 1.87, median of 0.82, and IQM of 0.94. These metrics demonstrate superior performance compared to existing model-based reinforcement learning baselines without look-ahead search techniques. For detailed quantitative analysis, Table 1 provides comprehensive scores for all games. The superior performance of EDELINE stems from its ability to preserve the visual fidelity advantages of diffusion-based world models. It architecture demonstrates significant improvements over DIAMOND in memory-intensive environments such as BankHeist and Hero. The enhanced performance arises from EDELINE’s memorization capabilities, combined with harmonizer-enabled precise visual detail selection for reward-relevant features. Appendices A, B, and C provide additional training curves, qualitative analyses, and performance metrics, respectively.

### v2

@@ 4 Motivational Experiments
(a) Qualitative results in MemoryS7
(b) Quantitative results in MemoryS7/S9
Figure 1: Motivational examples for both qualitative and quantitative evidences to demonstrate that DIAMOND face difficulties in imagining accurate future under memorization tasks.
To substantiate the memory limitations of the DIAMOND model, we conducted experiments using the MiniGrid MemoryS7 and MemoryS9 environments chevalier-boisvert2023minigrid . These experiments evaluate memory consistency and temporal prediction capabilities in world models. Fig. 1 presents qualitative comparisons among the DIAMOND model, our proposed EDELINE method, and the Oracle world model. The world models were trained on MiniGrid MemoryS7, and their state predictions were evaluated. The upper half of Fig. 1 (a) shows an initial action sequence of four consecutive left turns followed by four no-op actions. DIAMOND and EDELINE then autoregressively predicted the subsequent states. The lower half of Fig. 1 (a) reveals the prediction outcomes. At timestep 9, DIAMOND generated incorrect predictions with an extra key object in the state. The model’s predictions at timesteps 10 to 11 exhibited inaccuracies in the outer wall representations. These prediction errors result from the environment’s partial observability and DIAMOND’s four-frame context limitation. The input of four idle frames at timestep 9 led to loss of earlier state context. In contrast, EDELINE maintained accurate predictions throughout the sequence. This improved performance stems from EDELINE’s architectural design, which addresses the memory constraints inherent in DIAMOND. The accurate prediction at timestep 12 by DIAMOND can be attributed to its access to the 8th state, which enabled accurate predictions at timestep 13 through the correct 12th state context. Fig. 1 (b) illustrates DIAMOND’s performance limitations in both MemoryS7 and MemoryS9 environments due to insufficient memory capacity in partially observable scenarios. In contrast, EDELINE shows near-optimal performance under these conditions.
@@ 5 Methodology
Figure 2: The EDELINE world model includes three principal components: (1) A U-Net-like Next-Frame Predictor enhanced by adaptive group normalization and cross-attention mechanisms, (2) A Recurrent Embedding Module built on Mamba architecture for temporal sequence processing, and (3) A Reward/Termination Predictor implemented through linear layers. The EDELINE framework uses shared hidden representations across the components for efficient world model learning.
Conventional diffusion-based world models alonso2024diamond demonstrate promise in learning environment dynamics yet face fundamental limitations in memory capacity and horizon prediction consistency. To address these challenges, this paper presents EDELINE, as illustrated in Fig. 2, a unified architecture that integrates state space models (SSMs) with diffusion-based world models. EDELINE’s core innovation lies in its integration of SSMs for encoding sequential observations and actions into hidden embeddings, which a diffusion model then processes for future frame prediction. This hybrid design maintains temporal consistency while generating high-quality visual predictions. A Convolutional Neural Network based actor processes these predicted frames to determine actions, thus enabling autoregressive generation of imagined trajectories for policy optimization.
@@ 5.1.1 Recurrent Embedding Module
While DIAMOND, the current state-of-the-art in diffusion-based world models, relies on a fixed context window of four previous observations and actions sequence, the proposed EDELINE architecture advances beyond this limitation through a recurrent architecture for extended temporal sequence processing. At each timestep $t$, the Recurrent Embedding Module processes the current observation-action pair $(o_{t},a_{t})$ to update a hidden embedding $h_{t}=f_{\phi}(h_{t-1},o_{t},a_{t})$. The implementation of REM utilizes Mamba gu2024mamba , an SSM architecture that offers distinct advantages for world modeling. This architectural selection is motivated by the limitations of current sequence processing methods in deep learning. Self-attention-based Transformer architectures, despite their strong modeling capabilities, suffer from quadratic computational complexity which impairs efficiency. Traditional recurrent architectures including Long Short-Term Memory (LSTM) HochSchm97 and Gated Recurrent Unit (GRU) 69e088c8129341ac89810907fe6b1bfe experience gradient instability issues that affect dependency learning. In contrast, SSMs provide an effective alternative through linear-time sequence processing coupled with robust memorization capabilities via their state-space formulation. The adoption of Mamba emerges as a promising choice due to its demonstrated effectiveness in modeling temporal patterns across various sequence modeling tasks. Appendix D.9 presents a comprehensive ablation study that evaluates different architectural choices for the REM.
@@ 5.1.2 Next-Frame Predictor
While motivated by DIAMOND’s success in diffusion-based world modeling, EDELINE introduces significant architectural innovations in its Next-Frame Predictor to enhance temporal consistency and feature integration. At time step $t$, the model conditions on both the last $L$ frames and the hidden embedding $h_{t}$ from the Recurrent Embedding Module to predict the next frame $\hat{o}_{t+1}$. The predictive distribution $p_{\phi}(o^{0}_{t+1}|h_{t})$ is implemented through a denoising diffusion process, where $D_{\phi}$ functions as the denoising network. Let $y_{t}^{\tau}=(\tau,o^{0}_{t-L+1},...,o^{0}_{t},h_{t})$ represent the conditioning information, where $\tau$ represents the diffusion time. The denoising process can be formulated as $o^{0}_{t+1}=D_{\phi}(o^{\tau}_{t+1},y_{t}^{\tau}).$ To effectively integrate both visual and hidden information, $D_{\phi}$ employs two complementary conditioning mechanisms. First, the architecture incorporates Adaptive Group Normalization (AGN) AGN layers within each residual block to condition normalization parameters on the hidden embedding $h_{t}$ and diffusion time $\tau$, which establishes context-aware feature normalization AGN . This design significantly extends DIAMOND’s implementation, which limits AGN conditioning to $\tau$ and action embeddings only. The second key innovation introduces cross-attention blocks inspired by Latent Diffusion Models (LDMs), which utilize $h_{t}$ and $\tau$ as context vectors. The UNet’s feature maps generate the query, while $h_{t}$ and $\tau$ project to keys and values. This novel attention mechanism, which is absent in DIAMOND, facilitates the fusion of spatial-temporal features with abstract dynamics encoded in $h_{t}$. The observation modeling loss $\mathcal{L}_{\text{obs}}(\phi)$ is defined based on Eq. (9), and can be formulated as follows:
@@ 5.1.3 Reward / Termination Predictor
5.1.4 EDELINE World Model Training
To implement the loss functions defined in Eq. (2) and Eq. (3), we adopt a training strategy tailored to the architecture of the EDELINE world model. The Next-Frame Predictor is trained to reconstruct future observations from hidden embeddings sampled from arbitrary timesteps. In contrast to DIAMOND, which operates over a fixed four-frame window without memory, EDELINE maintains memory across the full trajectory. Directly applying the observation loss at every timestep, as done in DIAMOND, would incur significantly higher computational costs due to EDELINE’s recurrent embedding module. To mitigate this, we leverage the MAMBA parallel scan algorithm to efficiently compute hidden embeddings across all timesteps, and approximate the full loss by randomly sampling a single target timestep for reconstruction. Specifically, we sample a trajectory segment of fixed length $T$ from the replay buffer, indexed by =$\mathcal{I}=\{t,\ldots,t+T-1\}$. Given $(o^{0}_{i},a_{i})_{i\in\mathcal{I}}$, we initialize the hidden state $h_{t-1}$ and compute the hidden embeddings $\{h_{i}\}_{i\in\mathcal{I}}$ in parallel using MAMBA. We then randomly select a target timestep $j\in\{t+B,\ldots,t+T-1\}$, where $B$ denotes the burn-in period, and compute the observation reconstruction loss as:
$\mathcal{L}_{\text{obs}}(\phi)=\|\hat{o}^{0}_{j}-o_{j}\|^{2},\text{ where }j\sim\text{Uniform}\{t+B,\ldots,t+T-1\},\text{ and }\hat{o}^{0}_{j}\sim p(\hat{o}^{0}_{j}\mid h_{j-1}).$ (4)
This approach enables EDELINE to achieve computational efficiency comparable to DIAMOND, while retaining the advantages of long-term memory. For reward and termination prediction, EDELINE utilizes cross-entropy losses averaged over the sampled trajectory segment, which can be formulated as follows: $\mathcal{L}_{\text{rew}}(\phi)=\frac{1}{T}\sum_{i\in\mathcal{I}}\text{CrossEnt}(\hat{r}_{i},r_{i}),$ $\mathcal{L}_{\text{end}}(\phi)=\frac{1}{T}\sum_{i\in\mathcal{I}}\text{CrossEnt}(\hat{d}_{i},d_{i}).$ The detailed algorithm is presented in Appendix E. This unified training approach, combining random sampling strategies with dynamic loss harmonization, demonstrates superior efficiency compared to DIAMOND’s separate network methodology, as validated in our results presented in Section 6 and Appendix D.6. Moreover, the quantitative analysis presented in Appendix D.8 reveals substantial reductions in world model training duration.
The world model integrates an innovative end-to-end training strategy with a self-supervised approach. EDELINE extends the harmonization technique from HarmonyDream ma2024harmonydream through the adoption of harmonizers $w_{o}$ and $w_{r}$, which dynamically balance the observation modeling loss $\mathcal{L}_{\text{obs}}(\phi)$ and reward modeling loss $\mathcal{L}_{\text{rew}}(\phi)$. This adaptive mechanism results in the total loss function $\mathcal{L}(\phi)$:
$\begin{split}\mathcal{L}(\phi)=w_{0}\mathcal{L}_{\text{obs}}(\phi)&+w_{r}\mathcal{L}_{\text{rew}}(\phi)+\mathcal{L}_{\text{end}}(\phi)+\log(w_{o}^{-1})+\log(w_{r}^{-1}).\end{split}$ (5)
5.2 Agent Behavior Learning
To enable fair comparison and demonstrate the effectiveness of EDELINE’s world model architecture, the agent architecture adopts the same optimization framework as DIAMOND. Specifically, the agent integrates policy $\pi_{\theta}$ and value $V_{\theta}$ networks with REINFORCE value baseline and Bellman error optimization using $\lambda$-returns alonso2024diamond . The training framework executes a procedure with three key phases: experience collection, world model updates, and policy optimization. This method, as formalized in Algorithm 1, follows the established paradigms in model-based RL literature Kaiser2020SimPLe ; Hafner2020Dreamer ; micheli2023iris ; alonso2024diamond . To ensure reproducibility, we provide extensive details in the Appendix, with documentation of objective functions, and the hyperparameter configurations in Appendices F, D.12, respectively.
6 Experiments
This section presents our experimental results of EDELINE. We provide details of our setups in Appendix D.1. Additionally, we include ablation studies for each component of our approach in Appendix D.9. We also validate the benefits of our unified architecture through representation analysis in Appendix D.6, and evaluate the quality of our model’s imagination capabilities in Appendix D.7.
6.1 Atari 100k Experiments
@@
#Superhuman ($\uparrow$) 0 N/A 1 8 10 10 9 7 11 13
Mean ($\uparrow$) 0.000 1.000 0.332 0.956 1.046 1.266 1.124 0.989 1.459 1.866
Median ($\uparrow$) 0.000 1.000 0.134 0.505 0.289 0.580 0.485 0.270 0.373 0.817
IQM ($\uparrow$) 0.000 1.000 0.130 0.459 0.501 0.636 0.487 - 0.641 0.940
Optimality Gap ($\downarrow$) 1.000 0.000 0.729 0.513 0.512 0.433 0.510 - 0.480 0.387
@@
Following standard evaluation paradigms for world models, we evaluate EDELINE on the Atari 100k benchmark. For performance quantification, we adopt the human-normalized score (HNS) pmlr-v48-wangf16dueling , which measures agent performance relative to human and random baselines:
@@
HNS=agent scorerandom scorehuman scorerandom $\text{HNS}=\frac{\text{agent score}-\text{random score}}{\text{human score}-\text{random score}}$ (6)
Figure 3: Mean, Median and IQM HNS.
@@
Fig. 3 presents stratified bootstrap confidence intervals following agarwal2021deep ’s recommendations for point estimate limitations. It can be observed that EDELINE exhibits exceptional performance across this benchmark. Our approach surpasses human players in 13 games with a superhuman mean HNS of 1.87, median of 0.82, and IQM of 0.94. These metrics demonstrate superior performance compared to existing model-based reinforcement learning baselines without look-ahead search techniques. For detailed quantitative analysis, Table 1 provides comprehensive scores for all games. The superior performance of EDELINE stems from its ability to preserve the visual fidelity advantages of diffusion-based world models. It architecture demonstrates significant improvements over DIAMOND in memory-intensive environments such as BankHeist and Hero. The enhanced performance arises from EDELINE’s memorization capabilities, combined with harmonizer-enabled precise visual detail selection for reward-relevant features. Appendices D.3, D.4, and D.5 provide additional training curves, qualitative analyses, and performance metrics, respectively.
6.2 Crafter Experiments
Table 2: Comparison of different methods on Crafter in terms of average return and world model parameter count.
Method Avg Return #World Model Params
EDELINE 11.5 $\pm$ 0.9 11M
DreamerV3 XL 9.2 $\pm$ 0.3 200M
$\Delta$-IRIS 7.7 $\pm$ 0.5 25M
DreamerV3 M 6.2 $\pm$ 0.5 37M
IRIS 5.5 $\pm$ 0.7 48M
DIAMOND 2.8 $\pm$ 0.5 10.4M
To further evaluate EDELINE’s memory enhancement capabilities, we conducted experiments on Crafter hafner2022benchmarking , a procedurally generated survival environment that presents complex memory challenges. Crafter was specifically designed to assess ‘wide and deep exploration, long-term reasoning and credit assignment, and generalization’ hafner2024DreamerV3 , which establishes it as an ideal benchmark for the evaluation of an agent’s long-term memory utilization capabilities.
Crafter requires substantial memory capabilities due to its demand for agents to retain information about previously collected resources, crafted items, and explored territories for optimal decision-making, which establishes it as an ideal testbed for the evaluation of our memory-enhanced architecture. Within a 1M environment step budget, EDELINE achieves superior performance compared to state-of-the-art baselines including DIAMOND alonso2024diamond , DreamerV3 hafner2024DreamerV3 , $\Delta$-IRIS alonso2023delta-iris , and IRIS micheli2023iris , despite its relatively modest parameter count of 11M. These results highlight the significant advantages resulting from the integration of Mamba’s memory capabilities with diffusion’s generative abilities.
Table 2 presents our experimental results with a 1M environment step budget. The results reveal that EDELINE significantly outperforms all baselines with $25\%$ higher returns than DreamerV3 XL despite the utilization of 18$\times$ fewer parameters. Most importantly, EDELINE delivers a 4.1$\times$ improvement over DIAMOND with a comparable parameter count, which demonstrates the substantial benefits of our enhanced memory mechanism. In order to provide additional insight into EDELINE’s memory advantages, we conducted a qualitative analysis of model predictions over extended imagination. This analysis, as presented in Appendix D.11, visually demonstrates EDELINE’s superior ability to remember environmental features and maintain consistency when revisiting previously explored areas, a critical capability for effective planning in complex environments.